Suppose that the true linear regression model in a given situation is

7j9aVphhP4KfbsSEVO8mR4QAkZ+UWtVv7kYsV1vu

Now, assume that the researcher mistakenly believes that the true model is

ozUmrdo8Rjewz1VPZaeTswAAAABJRU5ErkJggg==,

and that he estimates this model, accordingly. Prove that his (OLS) estimator of iZsCXjnBFesNwU0897aTPv7Cb4QAAAABJRU5ErkJ will be biased.
Answer #1

LD LALA LULUS As trive Y . . ß + B x tu, is the mode & E (Y :) = 8 + Petri (E(ui) = 0] ELY) = pi + B, X. where 7 = IY & x = [.. p= Y is the O.L.S of fi for using the wrong model so, E($) = E(F)- E (1 TY :) n n i=1 Į (Bet patri) E (P) = 8 + f ₂ X . ß